PREFERRED STOCK RATE OF RETURN

What will be the nominal rate of return on a perpetual preferred stock with a $100 par value, a stated dividend of 11% of par, and a current market price of (a) $58.00, (b) $76.00, (c) $97.00, and (d) $145.00? Round your answers to two decimal places.

  1. %
  2. %
  3. %
  4. %
Answer #1

Annual dividend=100*11%=11

Nominal rate=Annual dividend/Current price

1.Nominal rate=(11/58)=18.97%(Approx).

2.Nominal rate=(11/76)=14.47%(Approx).

3.Nominal rate=(11/97)=11.34%(Approx).

4.Nominal rate=(11/145)=7.58%(Approx).